OpenStudy (anonymous):

log2 (-13.4) PLEASE HELP

OpenStudy (anonymous):

\[\log_{2}(-13.4) \]

OpenStudy (anonymous):

i only know how to take the ln of negative numbers based on euler's formula

OpenStudy (anonymous):

because \[e^{\pi*i} = -1\]

OpenStudy (anonymous):

@jim_thompson5910 can you help me please?

jimthompson5910 (jim_thompson5910):

let's see how far we get using that equation

jimthompson5910 (jim_thompson5910):

\[\Large e^{\pi*i} = -1\] \[\Large 13.4*e^{\pi*i} = 13.4*(-1)\] \[\Large 13.4*e^{\pi*i} = -13.4\] see where I'm going with this?

OpenStudy (anonymous):

okay this makes sense. but how will i take \[\log_{2} \] and clear e?

jimthompson5910 (jim_thompson5910):

by using the change of base formula \[\Large \log_{b}(x) = \frac{\ln(x)}{\ln(b)}\]

jimthompson5910 (jim_thompson5910):

the logs on the right side can be any base you want, in this case, they are base e

jimthompson5910 (jim_thompson5910):

just make sure that the logs on the right side have the same base

OpenStudy (anonymous):

so then would i say:\[\log_{2}(-13.4) = \frac{ \ln (-13.4) }{ \ln (2) } \] \[\frac{ \ln (-1) * \ln(13.4) }{ \ln(2) }\] \[\frac{ i \pi * \ln(13.4) }{ \ln(2) }\] and that would be my answer when i put the two natural logs in my calculator?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

but you use the rule log(x*y) = log(x)+log(y)

jimthompson5910 (jim_thompson5910):

that means ln(-1*13.4) = ln(-1) + ln(13.4) = pi*i + ln(13.4)

OpenStudy (anonymous):

oh okay thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome